A 60cm piece of wire is bent to form a right triangle. If the hypotenuse is 25cm, find the length of the longer of the two legs of the right triangle. (Show your solution)
Answer by mananth(16946)

A 60cm piece of wire is bent to form a right triangle. If the hypotenuse is 25cm, find the length of the longer of the two legs of the right triangle. (Show your solution)
...
perimeter = 60 cm
hypotenuse = 25
one leg = x
other leg = 60-(x+25)
=40+x
...
Pythagoras theorem
(35-x)^2+x^2=25^2
1225-70x+x^2+x^2=625
2x^2-70x+600=0
/2
x^2-35x+300=0
x^2-20x-15x+300=0
x(x-20)-15(x-20)=0
(x-20)(x-15)=0
x= 20 x=15
the longer leg = 20 cm
...
m.ananth@hotmail.ca